OpenStudy (anonymous):

y over negative 2 + 5= 13, what would y be

OpenStudy (anonymous):

Y / -2+5 = -4 -5 -5 Y / -2 = -9 *-2 *-2 Y = 18

OpenStudy (anonymous):

Is that good ?